OpenStudy (anonymous):

HELP! - prove that the function x sin 1/x is continuous for every value in the range of (0,1)

OpenStudy (anonymous):

using what? it is the composite of two continuous functions (well \(\frac{1}{x}\) is continuous on its domain) so it must be continuous

OpenStudy (anonymous):

using epsilon and delta.

OpenStudy (anonymous):

I think I have to prove it using for every d\[\delta >\ \left| x2-x1 \right| \] there exists also \[\epsilon > \left| f(x2)-f(x1) \right|\]

OpenStudy (anonymous):

Other way around. For every \(\varepsilon >0\) there exists a \(\delta >0\) so that \(|f(x)-f(x_0)|<\varepsilon\) for all \(x\) with \(|x-x_0|<\delta\). Think about it this way: \(\delta\) is a function of \(\varepsilon\).

OpenStudy (anonymous):

You should get someone else to help you, I'm sort of busy right now. Like this: @mukushla

OpenStudy (anonymous):

Yes, I really need help on this. I have ten questions that I need to solve by tomorrow or I will probably not pass on my course :(

OpenStudy (anonymous):

ok this is not so bad at all because the largest sine can be is 1

OpenStudy (anonymous):

let \(\epsilon>0\) be given and you want a \(\delta\) which you write in terms of \(\epsilon\) so that if \(|x-y|<\delta\) then \(x\sin(\frac{1}{x})-y\sin(\frac{1}{y})|<\epsilon\)

OpenStudy (anonymous):

last line is a typo should be \[|x\sin(\frac{1}{x})-y\sin(\frac{1}{y})|<\epsilon\]

OpenStudy (anonymous):

ok what's next? How do I get from here to the proof?

OpenStudy (anonymous):

oh sorry since the max of sine is 1, \[|x\sin(x)-y\sin(y)|\leq 2|x-y|\] and so you can pick \(\delta=\frac{\epsilon}{2}\)

OpenStudy (anonymous):

how did you get to x sin (x). Do you mean x sin (1/x)? Also how did you get to 2|x-y|? Where did the 2 come from? I'm just trying to understand it.